find the measures of the sides of triangle ABC and classify the triangle by its sides.
A (5,-2 ), B (7,2) C (3,5)

Not sur how to work this out.

Answer by stanbon(75887)

find the measures of the sides of triangle ABC and classify the triangle by its sides.
A (5,-2 ), B (7,2) C (3,5)
-----
AB: sqrt[(2--2)^2+(7-5)^2] = sqrt[16+4] = 2sqrt(5)
------------
AC: sqrt[(5--2)^2+(3--2)^2] = sqrt[(49)+25] = sqrt(74)
------------
BC: sqrt[(7-3)^2+(5-2)^2] = sqrt[16+9] = 5
--------------------------------------------
Triangle Type:: scalene
========================
Cheers,
Stan H.
========================